OpenStudy (anonymous):

Find S10 for the series 2 + 7 + 12 + 17 +...

OpenStudy (anonymous):

\[S_{10} = \frac n2[2a_1 + (n - 1)d]\]

OpenStudy (anonymous):

whats the anwser lol plz and thank you

OpenStudy (anonymous):

I'm sorry, but we're not allowed to just give you the answer...I can simplify it further though. \[S_{10} = \frac{10}{2}(2(2) + (10 - 1)5\]\[S_{10} = 5(4 + 45)\]\[S_{10} = 5(49)\]\[S_{10} = ?\]

OpenStudy (anonymous):

ohhh ok ty

OpenStudy (anonymous):

np :)